What number comes next in 3,9,27,81 number sequence
Found 2 solutions by Alan3354, TimothyLamb:
Answer by Alan3354(69443)

Get serious.
Make an effort.

Answer by TimothyLamb(4379)

3*3 = 9
3*9 = 27
3*27 = 81
answer:
3*81 = 243
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php